40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7114     Entry time: Wed Aug 8 10:15:13 2012
Author: jamie 
Type: Update 
Category: Environment 
Subject: Another earthquake, optics damped 

There were another couple of earthquakes at about 9:30am and 9:50am local.

earthquake.png

All but MC2 were off the watchdogs.  I damped and realigned everything and everything looks ok now.

Screenshot-Untitled_Window.png

ELOG V3.1.3-